Backlinks to DiaryRun9 in ARGOS Web (Search all webs)

Results from ARGOS web retrieved at 23:59 (GMT)

Commissioning
Commissioning run diaries Run 1st day Run ID Description 20111128 Run # 1 Nov 2011 Installation of dichroic interfaces 20130614 Run Installati...
DayTime20151005
SX T Gusset installation * At horizontal tube, turn upper shaft fully in. * At inclined tube, turn shaft only so far that there is the same distance in both...
DayTime20151006
Laser Tracker measurement on SX * Attachment of the retro plate on gusset. Extending the CalUnit holder focusing screw to avoid collision with DX. * Laser t...
DayTime20151007
Laser tracker measurement SX continue * Adjust of roll and tilt as well as position in park position with continuous check by laser tracker after each adjustme...
DayTime20151008
Rotation adjustment of CalUnit Rotation of CalUnit was adjusted within sub mm. The final position of the hexapod an swing arm for the final measurement have been:...
DayTime20151009
LAT testing The LAT was tested by AS in the electronic lab. End switch started to male function and didn't work anymore. It look like it something was with the te...
DayTime20151010
LAT mounting Electronic cabinet installed. LAT put on the telescope but could not be finally mounted. Upper taps for screws didn't fit. Lower ones are a bit to lo...
DayTime20151011
LAT Some adjustments of screws spacers and an additional bracket for the upper screws because they have been on the wrong place. Now the LAT is mounted. SX elect...
DayTime20151012
Swing arms Adjustment of working position on SX. Swap of park position switch to working position switch as this was confused before. Installation of park positio...
DayTime20151014
Horizon activities JZ starts to work on ?? up to 8:00. Zenith activities After telescope is back to Zenith we perform several task in parallel: i. Reconnect c...
DayTime20151016
LgswUnitDx configuration with IS DX LGSW DARK: 20151015_182131 We start redefining the subapertures: we modify the SX ones (20150330_170400). Pup center Subap...
DayTime20151017
It rained all night, even in the dome. 9:30: telescope back to Zenith we increase the air flow on both units: 4l/min on LGSW and 1.5l/min on PCD 16:00: CalUnit de...
DayTime20151018
10:00: startup and CalUnitSystemSx deployed. PockelsCellsTransmission on LgswUnitSx TN Plot Map ('20151018_164604','20151018_164724') Hexa ...
DayTime20151019
13:30: we try to calibrate the PatCam to Hexapod movement. We do steps pf 5" around the LGSW FoVs. SX LGSW PatCam to Hexapod IM: 20151019_203311 We determine also...
NightTime20151019
Laser Switch on SX at 16:25 local time. Power of ~45.4 Watts. SX Laser 1 has still its goose bump. Laser power measurement at 18:30 local time Laser 1: 16.057 Las...
NightTime20151020
16:00: power out up to now, we restart the system 19:00: TCS computer still down, we cannot move to Zenith... We exercise the AArb software, switching on and off ...
NightTime20151021
21:30: we switch on FLAO wfs sx, we ched the best focus for the pennina: 35mm SX Board 74.7 51.3 35 Then we check the best focus with the Hexapod, it ...
NightTime20151022
20:00: we start to prepare for opening the dome... 21:00: PockelsCellsTransmission on LgswUnitSx: TN Plot Map ('20151023_040836','20151023_040956') ...
NightTime20151023
17:30: Sx system started from AArb, we try a PockelsCellsTransmission test: TN Plot Map ('20151024_004135','20151024_004258') We start also the ...
Number of topics: 19
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback